OpenStudy (anonymous):

What is the equation of the ellipse centered at the origin with a vertex at (0,3) and co-vertex at (-2,0)?

OpenStudy (anonymous):

@BTaylor

OpenStudy (anonymous):

Here, this might explain it http://openstudy.com/study#/updates/512624cde4b0ce4f220669a9

OpenStudy (anonymous):

Kind of.. Could you help me out directly?

OpenStudy (anonymous):

Hmm...let me see

OpenStudy (anonymous):

what are the choices you have?

OpenStudy (anonymous):

A: \[\frac{ x^2 }{ 9 } + \frac{ y^2 }{ 4 } = 1\] They're all setup like that but the bottom numbers are different. B: x^2/3 + y^2/2 = 1 C: x^2/2 + y^2/3 = 1 D: x^2/4 + y^2/9 = 1

OpenStudy (anonymous):

I'm not good with this kind of thing